# CITRAS: Covariate-Informed Transformer for Time Series Forecasting

### v1

@@ Abstract
Covariates play an indispensable role in practical time series forecasting, offering rich context from the past and sometimes extending into the future. However, their availability varies depending on the scenario, and situations often involve multiple target variables simultaneously. Moreover, the cross-variate dependencies between them are multi-granular, with some covariates having a short-term impact on target variables and others showing long-term correlations. This heterogeneity and the intricate dependencies arising in covariate-informed forecasting present significant challenges to existing deep models. To address these issues, we propose CITRAS, a patch-based Transformer that flexibly leverages multiple targets and covariates covering both the past and the future forecasting horizon. While preserving the strong autoregressive capabilities of the canonical Transformer, CITRAS introduces two novel mechanisms in patch-wise cross-variate attention: Key-Value (KV) Shift and Attention Score Smoothing. KV Shift seamlessly incorporates future known covariates into the forecasting of target variables based on their concurrent dependencies. Additionally, Attention Score Smoothing transforms locally accurate patch-wise cross-variate dependencies into global variate-level dependencies by smoothing the past series of attention scores. Experimentally, CITRAS achieves state-of-the-art performance in both covariate-informed and multivariate forecasting, demonstrating its versatile ability to leverage cross-variate dependency for improved forecasting accuracy.
@@ 1 Introduction
Time series forecasting is a cornerstone in diverse fields such as energy [18, 35], retail [7, 5], and finance [6], where accurate predictions can drive strategic decisions and enhance operational efficiency. In practical scenarios, forecasters have access not only to the target variables they aim to predict but also to covariates that offer rich contexts for predictions. For instance, the electricity demand data shown in Figure 1 (Left) demonstrates a strong correlation with regional weather information and calendar data, highlighting the critical importance of appropriately incorporating covariates into the forecasting process.
@@
Driven by the pressing demands of real-world challenges, this paper addresses time series forecasting that includes two types of covariates. The first is the observed covariate, which has recorded values from the past up to the point of prediction. The second is the known covariate, which has values that are known from the past through to the future forecasting horizon, such as calendar events and promotion schedules.
@@
Figure 1: Left: Hourly electricity demand data for three weeks in Panama [2]. Temperature and humidity generally show a positive correlation with demand but exhibit local disturbances, as indicated by the red circle. Holiday and weekend indicators demonstrate a negative correlation with demand, providing important contextual information in future forecasting horizon. Right: Performance improvement of CITRAS through the utilization of past (observed and past known) covariates and future (future known) covariates.
@@
This practical problem setting presents two significant challenges. First, the forecasting model must be capable of universally accommodating heterogeneous variables. The availability and number of covariates can vary depending on the scenario, and situations often involve multiple target variables simultaneously, which is referred to as multivariate forecasting. Second, the model must capture the dependencies between variables from both fine-grained and coarse-grained perspectives. The fine-grained perspective enables the model to precisely capture the local impact of covariates, such as sudden drops in electricity demand on holidays. However, models that rely solely on a fine-grained perspective may overlook global dependencies between variables, as some covariates may be infrequent (e.g., holidays, promotions) or exhibit temporal disturbances. Therefore, it is crucial for the model to capture dependencies between variables at multiple levels to ensure robust and accurate forecasting.
@@
In recent years, deep forecasting models [33], particularly Transformers [32], have made remarkable progress. However, many studies have primarily focused on the dependencies among target variables. Some pioneering works [34, 24] have highlighted the importance of covariates; however, the effective integration of covariates, especially future known covariates, requires further exploration. Existing approaches to capture cross-variate dependencies can be broadly categorized into two types: variate-level and patch-level. iTransformer [22] is a representative variate-level approach that captures inter-variate correlations by embedding each sequence into a single variate token and applying attention across multiple variate tokens. On the other hand, patch-level approaches, represented by Crossformer [43], aim to capture more fine-grained cross-variate dependencies by applying attention between locally semantic tokens obtained through patching. However, these methods are polarizing, and an architecture capable of capturing both global and local dependencies has yet to be developed.
@@
To achieve effective time series forecasting that captures the intricate dependencies among diverse variables, we propose CITRAS: Covariate-Informed Transformer for time series forecasting. This model is a patch-based transformer that has a cross-time attention module and a cross-variate attention module separately. To effectively model the multi-granular cross-variate dependency, we introduce Attention Score Smoothing in the cross-variate attention module, which transforms locally accurate patch-wise dependencies into global variate-level dependencies. Furthermore, we introduce Key-Value (KV) Shift mechanism, which associates the key of the known covariate with the value from one patch step ahead, thereby seamlessly integrating future known covariates into the prediction process. These innovations expand the covariate understanding capabilities of the canonical Transformer, fully preserving its strong autoregressive properties.
@@
  • To fully utilize the rich context offered by covariates, we extend the Transformer to flexibly accommodate observed covariates and known covariates.
  • Technically, we introduce Attention Score Smoothing and KV Shift into the cross-variate attention module. These enhancements unravel the multi-granular cross-variate dependencies inherent in time series data and enable the effective utilization of future known covariates in predictions.
  • Our resulting model, CITRAS, achieves state-of-the-art performance on both covariate-informed settings and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependency for improved forecasting accuracy.
@@ 2.1 Transformer-based Time Series Forecasting
Given the significant success of Transformers in the fields of natural language processing [10] and computer vision [12], numerous studies have attempted to leverage their capabilities for time series forecasting. These works can be roughly divided into two approaches: channel-dependent (CD) and channel-independent (CI). The CD approach assumes that the future values of a variable are determined by its past values as well as the values of other variables. In contrast, the CI approach posits that a specific variable only depends on its past values, omitting interactions between variables.
Early applications of Transformers in time series forecasting can be viewed as CD approaches, where multiple variables at the same time point are embedded into temporal tokens [44, 40, 45, 20, 11]. To address the computational costs and intricate dependencies arising from long time series data, Informer [44] designs a ProbSparse self-attention mechanism with a distilling operation to efficiently focus on important features. Autoformer [40] incorporates decomposition and an auto-correlation mechanism to uncover reliable dependencies from complex temporal patterns. FedFormer [45] employs frequency-domain sparse representations to efficiently capture the global properties of time series. However, point-wise temporal tokens have limited local semantics, making it difficult to capture intricate cross-time and cross-variate dependencies.
@@
However, the CI approach is sometimes regarded as an oversimplification [1], especially considering that fluctuations in time series data are often influenced by external factors. A representative CD approach is iTransformer [22], which embeds each channel into a single variate token and applies cross-variate attention over them. TimeXer [34] also captures cross-variate dependencies at the variate level, with the additional use of cross-time attention to capture intra-variate temporal dependencies. Meanwhile, Crossformer [43] applies cross-variate attention to patch-level representations. Recent Any-variate Attention techniques [37, 24, 21] also apply attention to patch-level representations obtained by flattening all variables into one sequence. While the patch-level approach captures more fine-grained dependencies between variables, its local receptive field lacks global context awareness. Our method, CITRAS, first captures fine-grained dependencies at the patch level, and Attention Score Smoothing gradually refines them to achieve global variate-level dependencies.
@@ 2.2 Covariate-Informed Time Series Forecasting
To leverage the rich context offered by covariates, statistical models such as ARIMAX [36] and VARX [26] extend the original models by assuming linear relationships between variables. To capture more intricate dependencies through deep learning, N-BEATSx [28] enhances the original N-BEATS [29] by introducing an additional exogenous stack. TiDE [8] proposes an MLP-based architecture that utilizes future known covariates in the final projection to leverage their direct effects on future target variables. Temporal Fusion Transformer [19] designs gating mechanisms on top of the Transformer to select relevant features for forecasting. While these approaches adopt elaborate model designs to capture covariates, recent methods, such as TimeXer [34] and Timer-XL [24], aim to integrate covariates by leveraging the inherent expressive power of the Transformer. We follow this line of work and extend the Transformer to accommodate covariates while fully preserving the original architecture. Additionally, we focus primarily on the utilization of future known covariates, which, despite their importance, have not been sufficiently discussed in these studies.
@@ 3.2 Architecture
First, all sequences are segmented into non-overlapping patches of length $P$, and each patch is embedded into a $D$-dimensional token. Taking target variables as an example, this can be formalized as:
@@
where $N_{tgt}=\lfloor\frac{T}{P}\rfloor$ is the number of patches and $\operatorname{Embed}:\mathbb{R}^{P}\to\mathbb{R}^{D}$ is a shared linear projector across all variables. We denote the token embeddings of a target variable at all patch steps as $\mathbf{H}_{:}^{tgt,c}=~{}\{\mathbf{H}_{i}^{tgt,c}\}_{i=1}^{N_{tgt}}~{}\in~{}\mathbb{R}^{N_{tgt}\times D}$. Similarly, the token embeddings of an observed covariate are denoted as $\mathbf{H}_{:}^{obs,c}\in\mathbb{R}^{N_{obs}\times D}$, and those of a known covariate are denoted as $\mathbf{H}_{:}^{knw,c}\in\mathbb{R}^{N_{knw}\times D}$, all obtained using the shared parameters with the target variables. Here, $N_{obs}=N_{tgt}=\lfloor\frac{T}{P}\rfloor$ and $N_{knw}=\lfloor\frac{(T+S)}{P}\rfloor$.
@@
where $\operatorname{LN}$ denotes layer normalization [4], $\operatorname{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ denotes the multi-head attention layer where $\mathbf{Q}$,$\mathbf{K}$, and $\mathbf{V}$ serve as queries, keys and values, and $\operatorname{FFN}$ denotes a feed-forward network.
Again, token embeddings of observed covariates $\mathbf{H}_{:}^{obs,c}$ and those of known covariates $\mathbf{H}_{:}^{knw,c}$ are similarly processed with the shared parameters.
@@
where $\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{D\times d_{k}}$ and $d_{k}$ is the dimension of the query, key, and value. In the cross-variate attention module, this attention score is calculated for each patch step separately, capturing fine-grained dependency between variables at that step. However, time series data often exhibit local disturbances, and sparse variables may maintain constant values within a patch. In such scenarios, local patch-wise attention fails to represent global variate-level dependencies.
@@
where $\mathbf{A}_{1}=\widetilde{\mathbf{A}}_{1}$ and $\alpha$ is a smoothing factor, which we consider as a shared hyperparameter across all cross-variate attention modules and heads. Note that the aim and structure of Attention Score Smoothing are completely different from the exponential smoothing attention in ETSformer [38], which adopts an exponentially decreasing weights based on the time lag instead of the vanilla attention scores to capture intra-variate cross-time dependency.
@@
Following the next token prediction approach common in decoder-only Transformers, each token embedding of the target variables is used to predict the values of the next patch. For $i=1,...,N_{tgt}$ and $c=1,...,C_{tgt}$, this can be formalized as:
@@
where $\operatorname{Project}:\mathbb{R}^{D}\to\mathbb{R}^{P}$ is a shared linear projector across all steps and target variables. In the training phase, we calculate the squared loss using all of these outputs.
@@ 4 Experiments
To verify the effectiveness and versatility of CITRAS, we extensively evaluated it in the two settings, short-term covariate-informed forecasting and well-acknowledged long-term multivariate forecasting.
@@ 4.1 Covariate-Informed Forecasting
For covariate-informed forecasting, we use the short-term electricity price forecasting (EPF) dataset [18] and electricity demand forecasting (EDF) dataset [2]. The EPF dataset consists of six years of hourly time series data from five major power markets, with electricity price as the target and day-ahead load forecasts and generation forecasts as known covariates. The EDF dataset is an hourly time series dataset spanning over five years, where the target variable is the national electricity demand in Panama. We adopt four meteorological indices measured in three cities (total 12 indices) as observed covariates, and three binary calendar indicators (holiday, weekend, school day) as known covariates. Following previous works [28], we set the input length as 168, forecasting length as 24, and patch length as 24. Further details about the datasets are included in the Appendix A.1.
@@
We include eight strong deep forecasting models, namely Timer-XL[24], TimeXer [34], iTransformer [22], TiDE [8], Temporal Fusion Transformer (TFT) [19], Crossformer [43], TimesNet [39], and Autoformer [40]. Notably, TiDE and TFT are originally developed to accommodate future known covariates. For Timer-XL, TimeXer and iTransformer, we extend their original implementation to support future known covariate, as detailed in the Appendix A.3. Crossformer, TimesNet and Autoformer remain original implementation and cannot accommodate future known covariates.
@@
Comprehensive results for covariate-informed forecasting are presented in Table 1. In the EPF dataset, the day-ahead load forecasts and generation forecasts show a strong correlation with future electricity prices, providing valuable insights for forecasting. Crossformer, TimesNet, and Autoformer lack the ability to capture these known covariates, resulting in inferior performance. iTransformer and TimeXer capture these covariates at the variate level, while Timer-XL captures them at the patch level. These contrasting approaches show competitive performance. On the other hand, CITRAS incorporates these covariates at a locally accurate patch level and captures dependencies at the variate level as well, demonstrating superior accuracy. The EDF dataset poses further challenges to the models due to the presence of numerous observed and known covariates, some of which exhibit local disturbances. Interestingly, Timer-XL, which applies attention to all token embeddings at once, performs worse than other variate-level approaches due to noisy interactions. In contrast, CITRAS applies separate attention to cross-time and cross-variate directions and employs Attention Score Smoothing, allowing it to robustly capture dependencies despite disturbances, thereby outperforming other baselines.
Table 1: Full results of the covariate-informed forecasting, where future known covariates are available. We adhere to the standard protocol [28], setting the input length and forecasting length to 168 and 24, respectively. Avg represents the average results from all EPF datasets. EPF results for Crossformer, TimesNet, and Autoformer are cited from [24], while others are reproduced. An asterisk (*) indicates that the model cannot accommodate future known covariates.
@@
Cross.*
TimesNet*
Auto.*

### v2

@@ Abstract
In practical time series forecasting, covariates provide rich contextual information that can potentially enhance the forecast of target variables. Although some covariates extend into the future forecasting horizon (e.g., calendar events, discount schedules), most multivariate models fail to leverage this pivotal insight due to the length discrepancy with target variables. Additionally, capturing the dependency between target variables and covariates is non-trivial, as models must precisely reflect the local impact of covariates while also capturing global cross-variate dependencies. To overcome these challenges, we propose CITRAS, a decoder-only Transformer that flexibly leverages multiple targets, past covariates, and future covariates. While preserving strong autoregressive capabilities, CITRAS introduces two novel mechanisms in patch-wise cross-variate attention: Key-Value (KV) Shift and Attention Score Smoothing. KV Shift seamlessly incorporates future covariates into the forecasting of target variables based on their concurrent dependencies. Additionally, Attention Score Smoothing refines locally accurate patch-wise cross-variate dependencies into global variate-level dependencies by smoothing the past series of attention scores. Experimentally, CITRAS outperforms state-of-the-art models on thirteen real-world benchmarks from both covariate-informed and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependencies for improved forecasting accuracy.
@@ 1 Introduction
Time series forecasting is a cornerstone in diverse fields such as energy [23, 44], retail [9, 6], and finance [7], where accurate predictions can drive strategic decisions and enhance operational efficiency. In practical scenarios, forecasters have access not only to the target variables they aim to predict but also to covariates that offer rich contexts for predictions. For instance, the "Electricity Demand" data shown in Figure 1 (Left) demonstrates a strong negative correlation with the "Holiday" indicator, which is accessible up to the future forecasting horizon. This highlights the critical importance of appropriately incorporating covariates into the forecasting process.
@@
Driven by these practical demands, this paper addresses time series forecasting that includes two types of covariates. The first is the observed covariate, which has recorded values from the past up to the point of prediction. The second is the known covariate, which has values available from the past through to the future forecasting horizon. Known covariates can take predetermined (e.g., calendar events), estimated (e.g., weather forecasts), or controllable values (e.g., discounts in retail). For clarity, we sometimes refer to the past values of observed covariates and known covariates collectively as past covariates, and the future values of known covariates as future covariates.
@@
Figure 1: Left: Hourly electricity demand data for three weeks in Panama [2]. "Holiday" and "Weekend" indicators demonstrate a negative correlation with demand, providing important contextual information in the future forecasting horizon. "Humidity" displays a weak correlation but introduces temporal disturbances (red circle), which may distract the forecasting process. Right: Performance improvement of CITRAS through the utilization of past covariates and future covariates.
@@
This practical problem setting presents two significant challenges. First, the forecasting model must be capable of universally accommodating heterogeneous variables. The availability and number of variables can vary depending on the scenario, and known covariates exhibit a length discrepancy compared to other variables. Second, the model must capture the dependencies between variables from both fine-grained and coarse-grained perspectives. The fine-grained perspective enables the model to precisely capture the local impact of covariates, such as sudden drops in electricity demand on holidays. However the model that relies solely on a fine-grained perspective may fail to capture global dependencies between variables, as it can overlook covariates representing infrequent events or can be distracted by temporal disturbances. Therefore, it is crucial for the model to capture dependencies between variables at both fine-grained and coarse-grained levels to ensure robust and accurate forecasting.
@@
In recent years, deep forecasting models [42], particularly Transformers [40], have made remarkable progress. However, most Transformer-based models have primarily focused on multivariate settings that consider only multiple target variables. Some pioneering works [43, 30] have highlighted the importance of covariates; yet, the effective integration of heterogeneous covariates that leverages the inherent expressive power of Transformers requires further exploration. Existing approaches to capture cross-variate dependencies can be broadly categorized into two types: variate-level and patch-level. iTransformer [28] is a representative variate-level approach that captures inter-variate correlations by embedding each sequence into a single variate token and applying attention across multiple variate tokens. On the other hand, patch-level approaches, represented by Crossformer [55], applies cross-variate attention between locally semantic tokens obtained through patching. While variate-level approaches excel at capturing global dependency robustly against noisy interaction, patch-level approaches are better at capturing fine-grained information. Unfortunately, an architecture that enjoys the advantages of both of these approaches is yet to be developed.
@@
To achieve effective time series forecasting that captures the intricate dependencies among diverse variables, we propose CITRAS: Covariate-Informed Transformer for time series forecasting. This model is a decoder-only patch-based Transformer that has a cross-time attention module and a cross-variate attention module separately. To effectively model the cross-variate dependency at both fine-grained and coarse-grained perspectives, we introduce Attention Score Smoothing in the cross-variate attention module, which refines locally accurate patch-wise dependencies into global variate-level dependencies. Furthermore, we introduce Key-Value (KV) Shift mechanism, which associates the key of the known covariate with the value from one patch step ahead, thereby seamlessly integrating future covariates into the prediction process. These innovations expand the covariate understanding capabilities of the canonical decoder-only Transformer, fully preserving its strong autoregressive properties.
@@
  • To fully utilize the rich context offered by covariates, we extend the decoder-only Transformer to flexibly accommodate observed covariates and known covariates while fully maintaining its original strong autoregressive capability.
  • Our model, CITRAS, introduces two novel mechanisms into the cross-variate attention module. Attention Score Smoothing captures locally accurate patch-wise dependencies and refines them into global variate-level dependencies, while KV Shift seamlessly integrates future covariates based on the obtained dependencies.
  • Experimentally, CITRAS outperforms state-of-the-art models on thirteen benchmarks from both covariate-informed and multivariate settings, demonstrating its versatile ability to leverage cross-variate and cross-time dependencies for improved forecasting accuracy.
@@ 2.1 Transformer-based Time Series Forecasting
Given the significant success of Transformers in the fields of natural language processing [12] and computer vision [15], numerous studies have attempted to leverage their capabilities for time series forecasting. These works can be roughly divided into two approaches: channel-dependent (CD) and channel-independent (CI). The CD approach assumes that the future values of a variable are determined by its past values as well as the values of other variables. In contrast, the CI approach posits that a specific variable only depends on its past values, omitting explicit interactions between variables.
Early applications of Transformers in time series forecasting can be viewed as CD approaches, where multiple variables at the same time point are embedded into temporal tokens [56, 49, 57, 26, 13]. To address the computational costs and intricate dependencies arising from long time series data, Informer [56] designs a ProbSparse self-attention mechanism with a distilling operation to efficiently focus on important features. Autoformer [49] incorporates decomposition and an auto-correlation mechanism to uncover reliable dependencies from complex temporal patterns. However, point-wise temporal tokens have limited local semantics, making it difficult to capture intricate cross-time and cross-variate dependencies.
@@
However, the CI approach is sometimes regarded as an oversimplification [1], especially considering that fluctuations in time series data are often influenced by external factors. A representative CD approach is iTransformer [28], which embeds each channel into a single variate token and applies cross-variate attention over them. TimeXer [43] and Leddam [51] also capture cross-variate dependencies at the variate level, with the additional cross-time attention to capture intra-variate temporal dependencies. Meanwhile, Crossformer [55] applies cross-variate attention to patch-level representations. Any-variate Attention techniques [46, 30, 27] also apply attention to patch-level representations obtained by flattening all variables into one sequence. While the patch-level approach captures more fine-grained dependencies between variables, its local receptive field lacks global context awareness. Our method, CITRAS, first captures fine-grained dependencies at the patch level, and Attention Score Smoothing gradually refines them to achieve global variate-level dependencies.
@@ 2.2 Covariate-Informed Time Series Forecasting
To leverage the rich context offered by covariates, numerous approaches have been proposed, including statistical models [45, 33], MLP variants [10, 35, 8, 16], RNN variants [38], and Transformer-based models [24]. For instance, TiDE [10] introduces an MLP-based architecture that utilizes future covariates in the final projection to leverage their direct effects on future target variables. TSMixer-Ext [8] embeds past and future covariates into the same shape and applies feature mixing in the early stage. Temporal Fusion Transformer [24] designs gating mechanisms on top of the Transformer to select relevant features for forecasting. Contemporary to our work, ChronosX [4] introduces injection blocks to integrate covariates into pretrained models. A common drawback of these models is their structural complexity, which arises from architectural additions needed to handle heterogeneous variables.
Recent works, such as TimeXer [43] and Timer-XL [30], have attempted simpler integration of covariates into Transformers to retain their inherent expressive power. However, these approaches require further exploration. Specifically, the variate-level approach adopted by TimeXer is not compatible with decoder-only Transformers, leaving room for enhanced autoregressive capabilities through next-token prediction. The masking scheme in Timer-XL, known as TimeAttention, is not designed to accommodate future covariates, raising questions about its ability to utilize these pivotal features. In contrast to these works, CITRAS seamlessly integrates both observed and known covariates into the decoder-only Transformer, fully preserving its strong autoregressive capability.
@@ 3.2 Architecture
First, all sequences are segmented into non-overlapping patches of length $P$, and each patch is embedded into a $D$-dimensional token. For simplicity, we assume that both $T$ and $S$ are divisible by $P$. Taking target variables as an example, this can be formalized as:
@@
where $N_{tgt}=\frac{T}{P}$ is the number of patches and $\operatorname{Embed}:\mathbb{R}^{P}\to\mathbb{R}^{D}$ is a shared linear projector across all variables. We denote the token embeddings of a target variable at all patch steps as $\mathbf{H}_{:}^{tgt,c}=~{}\{\mathbf{H}_{i}^{tgt,c}\}_{i=1}^{N_{tgt}}~{}\in~{}\mathbb{R}^{N_{tgt}\times D}$. Similarly, the token embeddings of an observed covariate are denoted as $\mathbf{H}_{:}^{obs,c}\in\mathbb{R}^{N_{obs}\times D}$, and those of a known covariate are denoted as $\mathbf{H}_{:}^{knw,c}\in\mathbb{R}^{N_{knw}\times D}$, all obtained using the shared parameters with the target variables. Here, $N_{obs}=N_{tgt}=\frac{T}{P}$ and $N_{knw}=\frac{(T+S)}{P}$.
@@
where $\operatorname{LN}$ denotes layer normalization [5], $\operatorname{MHA}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)$ denotes the multi-head attention layer where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ serve as queries, keys and values, and $\operatorname{FFN}$ denotes a feed-forward network. Again, token embeddings of observed covariates $\mathbf{H}_{:}^{obs,c}$ and those of known covariates $\mathbf{H}_{:}^{knw,c}$ are similarly processed with the shared parameters.
@@
where $\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{D\times d_{k}}$ and $d_{k}$ is the dimension of the query, key, and value. In the cross-variate attention module, this attention score is calculated for each patch step separately, capturing fine-grained dependency between variables at that step. However, time series data often exhibit local disturbances, and sparse variables representing infrequent events may maintain constant values within a patch. In such scenarios, local patch-wise attention fails to represent global variate-level dependencies.
@@
where $\mathbf{A}_{1}=\widetilde{\mathbf{A}}_{1}$ and $\alpha$ is a smoothing factor, which we consider as a shared hyperparameter across all cross-variate attention modules and heads. It is important to note that the aim and structure of Attention Score Smoothing are completely different from existing integrations of EMA into attention mechanisms, as attempted in ETSformer [47] or CARD [41]. These works integrate EMA into cross-time attention to robustly capture intra-variate variations. In contrast, we apply EMA to cross-variate attention to refine cross-variate dependencies.
@@
Following the next-token prediction approach common in decoder-only Transformers, each token embedding of the target variables is used to predict the values of the next patch. For $i=1,...,N_{tgt}$ and $c=1,...,C_{tgt}$, this can be formalized as:
@@
where $\operatorname{Project}:\mathbb{R}^{D}\to\mathbb{R}^{P}$ is a shared linear projector across all steps and target variables. In the training phase, we calculate the squared loss using all of these outputs. In the testing phase, the outputs from the last patches are used for forecasting. When the forecasting horizon $S$ exceeds the patch length $P$, the output target values are integrated into subsequent input for recursive forecasting.
@@ 4 Experiments
To verify the effectiveness and versatility of CITRAS, we extensively evaluated it in the two settings, covariate-informed forecasting and well-acknowledged multivariate forecasting.
@@ 4.1 Covariate-Informed Forecasting
We use seven real world datasets for covariate-informed forecasting. (1)-(5) EPF [23]: Electricity Price Forecasting datasets of five major power markets spanning six years. The target variable is the hourly electricity price and we adopt day-ahead load forecasts and generation forecasts as known covariates. (6) EDF [2]: Electricity Demand Forecasting dataset in Panama spanning over five years. The target variable is the hourly national electricity demand in Panama. We adopt four meteorological indicators measured in three cities (total 12 indices) as observed covariates, and three binary calendar features (holiday, weekend, school day) as known covariates. (7) BS [17]: Bike Sharing dataset recording the hourly rental counts of Bikeshare system in Washington D.C. spanning two years. The target variables are three types of rental counts. We adopt five meteorological indicators as observed covariates and three calendar features (holiday, weekday, working day) as known covariates. Following previous works [35], we set the input length as 168 and forecasting length as 24. Further details about the datasets are included in Appendix A.1.
@@
We include the following baselines, categorized into four groups. (A) CD models that originally support future covariates, including TSMixer-Ext [8], TiDE [10], Temporal Fusion Transformer (TFT) [24], and TimeXer [43]. (B) CD models that are extended from original designs (as detailed in A.3) to support future covariates, including Timer-XL [30], iTransformer [28], and Leddam [51]. (C) CD models that cannot accommodate future covariates, including CARD [41], ModernTCN [14], TimesNet [48], and Crossformer [55]. (D) CI models including FITS [50], DLinear [52], and PatchTST [34]. For fair comparison with baselines (C) and (D), we also evaluate CITRAS without future covariates (i.e., without using KV Shift).
@@
Comprehensive results for covariate-informed forecasting are presented in Table 1. We observe that recent models capable of utilizing future covariates outperform those that cannot, underscoring the indispensability of future covariates in practical forecasting scenarios. Among these models, TimeXer captures cross-variate dependency at the variate level, while Timer-XL captures it at the patch level. The superiority of these contrasting methods is dataset-dependent. In contrast, CITRAS leverages the benefits of both approaches and seamlessly integrates future covariates, consistently demonstrating strong performance. Moreover, the superiority of CITRAS remains evident even when evaluated using conventional protocols that do not employ future covariates. This is because ordinary multivariate models treat target variables and covariates equally, leading to unnecessary interaction between covariates. In contrast, CITRAS applies attention only from target variables to covariates, avoiding this negative interaction.
Table 1: Full results of the covariate-informed forecasting with and without future covariates. The input length and forecasting length is set to 168 and 24, respectively. EPF results for "w/o future" are cited from TimeXer  [30] if available; other results are reproduced.
@@
TimeXer
Timer-XL
iTrans.
Leddam
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
w/ future
EPF–NP
0.171
0.214
0.187
0.253
0.296
0.309
0.193
0.237
0.175
0.225
0.176
0.220
0.203
0.250
0.227
0.266
EPF-PJM
0.063
0.153
0.078
0.174
0.105
0.213
0.079
0.174
0.080
0.175
0.062
0.155
0.077
0.171
0.087
0.193
EPF–BE
0.362
0.248
0.316
0.241
0.447
0.301
0.352
0.240
0.368
0.231
0.384
0.257
0.339
0.246
0.369
0.253
EPF–FR
0.363
0.176
0.396
0.214
0.407
0.251
0.371
0.182 0.364
0.193
0.430
0.214
0.378
0.205
0.416
0.221
EPF–DE
0.217
0.293
0.250
0.322
0.515
0.462
0.262
0.326
0.283
0.322
0.228 0.299
0.258
0.327
0.273
0.332
EDF
0.070
0.193
0.077
0.209
0.133
0.259
0.092
0.225
0.071
0.199
0.079
0.206
0.071 0.197
0.080
0.206
BS
0.280
0.313
0.354
0.363
0.488
0.446
0.318
0.346
0.338
0.357
0.286 0.318
0.358
0.360
0.376
0.384
Model
CITRAS
CARD
ModernTCN
TimesNet
Cross.
FITS
DLinear
PatchTST
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
w/o future
EPF–NP
0.226
0.259
0.256
0.287
0.234 0.272
0.250
0.289
0.240
0.285
0.304
0.316
0.309
0.321
0.267
0.284
EPF-PJM
0.088
0.181
0.108
0.211
0.101
0.195 0.097 0.195
0.101
0.199
0.109
0.216
0.108
0.215
0.106
0.209
EPF–BE
0.405
0.275
0.414
0.278
0.383
0.260
0.419
0.288
0.420
0.290
0.460
0.308
0.463
0.313
0.400 0.262
EPF–FR
0.408
0.220
0.420
0.243
0.407
0.224
0.431
0.234
0.434
0.208
0.402
0.254
0.429
0.260
0.411
0.220
EPF–DE
0.408
0.400
0.453
0.424
0.444 0.416
0.502
0.446
0.574
0.430
0.533
0.473
0.520
0.463
0.461
0.432
EDF
0.086
0.205
0.088
0.211
0.086
0.214
0.101
0.233
0.108
0.233
0.134
0.258
0.129
0.252
0.103
0.232
BS
0.289
0.320
0.416
0.414
0.378
0.402
0.363
0.366
0.327
0.354
0.484
0.439
0.469
0.429
0.378
0.348
4.2 Multivariate Forecasting
For a comprehensive comparison, we further evaluate our model on well-recognized multivariate forecasting datasets without using any covariates. To reflect practical scenarios, we adopt the rolling forecasting approach [29], where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works [29, 30], we set the input length to 672, the training forecasting length to 96, and the testing forecasting length to {96, 192, 336, 720}. We present the average results for the four forecasting lengths in Table 2. Compared to recent models specifically designed for these multivariate settings, CITRAS achieves superior or competitive performances in most datasets, highlighting the effectiveness and versatility of its design across various settings.
Table 2: Results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. We adopt the rolling forecasting approach  [29], where one model is used for four forecasting lengths. Averaged results are reported here and full results are provided in Table 10. Results are cited from Timer-XL [30] if available; otherwise reproduced.
Model
CITRAS
TimeXer
Timer-XL
iTrans.
Leddam
CARD
ModernTCN
FITS
DLinear